OpenStudy (anonymous):

Need help with a pre-calculus question I can't get the statements to be true. Write statements S1, S2 and S3. Sn: 2+5+8+...+(3n-1)=n(1+3n/2). For S1 I get 5=7. For S2, I get 14=40. And S3 23=100. How do I make these true statements?

OpenStudy (anonymous):

well.....

OpenStudy (anonymous):

Hi.

OpenStudy (anonymous):

give me a sec

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

i think pax has it

OpenStudy (anonymous):

lol its been a while since ive taken this

OpenStudy (anonymous):

how do I find pax?

OpenStudy (paxpolaris):

\[S_n: 2+5+8+...+(3n-1)={n(1+3n)\over 2}\]

OpenStudy (anonymous):

So i go (3(2)-1)=2(1+3(2)/2 = -5=7 so it is false. How do I make this statement true?

OpenStudy (anonymous):

Sorry, I meant = 5=7

OpenStudy (paxpolaris):

\[S_1: 2={1(1+3\cdot1)\over 2}\]\[S_2: 2+5={2(1+3\cdot2)\over 2}\]\[S_3: 2+5+8={3(1+3\cdot3)\over 2}\]

OpenStudy (anonymous):

I got it. 2 is the first and we add 1 to the other side. In S2, we add 2+5 and 2 to the other side. S3, we add the 2+5+8 and on the other side, we add 3. It is the progression of the numbers. Thank you.

OpenStudy (paxpolaris):

right :) S3 means the sum of the first 3 numbers of the sequence and in the right you have the formula for finding the sum using n=3

OpenStudy (anonymous):

Perfect. I got S1: 2=2. S2: 7=7 and S3: 15=15 Thank you Pax. I appreciate the help.